show that the set {(1,2,3),(0,1,2),(0,0,1) of vectors generates or span R^3


1
Expert's answer
2021-06-15T04:55:36-0400

We can set up a matrix and use Gaussian elimination to figure out the dimension of the space they span.


"\begin{pmatrix}
 1 & 2 & 3 \\
 0 & 1 & 2 \\
 0 & 0 & 1 \\
\end{pmatrix}"

"R_1=R_1-2R_2"


"\begin{pmatrix}
 1 & 0 & -1 \\
 0 & 1 & 2 \\
 0 & 0 & 1 \\
\end{pmatrix}"

"R_1=R_1+R_3"


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 2 \\
 0 & 0 & 1 \\
\end{pmatrix}"

"R_2=R_2-2R_3"


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 1 \\
\end{pmatrix}"

Since the rank of the matrix is 3, then the set {(1,2,3),(0,1,2),(0,0,1) of vectors generates or span "R^3."